Show that:  T=I/Y


1
Expert's answer
2021-06-29T13:08:10-0400

Show that

"T=\frac{l}{Y}"

We know that

Normal stress="\frac{mg}{\pi r^2}"

Longitudinal strain="\frac{l}{L}"

Young modules constant =(normal stress /longitudinal strain)

"Y=\frac{mgL}{\pi r^2l}\rightarrow(1)"

Where

"T=mg"

Let us assume wire lenth L=1 m and area(A)= 1m2 then we can written as

"\frac{L}{\pi r^2}=1m"

Put value equation (1)

Then we can written as

"Y=\frac{l}{T}"

"T=\frac{l}{Y}"

Hence proved